# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ 5 Experiments
In this section, we conduct experiments to understand equilibrium outcomes and, more generally, how the allocation parameter $\rho$ impacts creator behavior. We consider cost functions of the form $c_{i}(x)=a_{i}x^{\theta}$. In the default configuration, the parameters are $n=20$, $\gamma=0.8$, $\mu=100$, $\theta=1.5$, $\alpha=0.5$, and the coefficients $\{a_{i}\}_{i=1}^{n}$ are independently drawn from the uniform distribution $Uni[1,10]$. In addition to the default configuration, we vary one parameter at a time while keeping all others fixed at their default values to assess its individual impact. Each point in Figure 2 represents the average outcome across 150 independent game instances, with error bars indicating 95% confidence intervals obtained via bootstrapping. We discretize the possible values of $\rho$ into 100 equally spaced points in the range $[0,1]$. For each $\rho$, we check if it induces an $\varepsilon$-FSE in $G(\rho)$ for $\varepsilon=10^{-4}$. We ran all the executions on a standard laptop, completing the entire process in a few hours. We include all implementation details in Appendix D.
@@
Minimal revenue-allocation threshold.
Figure 2(a) illustrates how the minimal allocation parameter $\rho$ required for inducing an FSE varies with ecosystem parameters. All curves decrease with the number of creators, as Theorem 5 implies: The content of each creator becomes insignificant w.r.t. all other created content, so withholding has a high alternative cost. Furthermore, more powerful GenAI (larger $\alpha$) enables lower $\rho$ choices for stable FSE, as the curves associated with a change in $\alpha$ (red, green, orange, and black curves) are ordered. The black curve $(\mu=100)$ lies above the blue curve $(\mu=50)$, suggesting that creators demand more compensation as traffic increases.
Content quality vs allocation parameter
Figure 2(b) shows how different values of $\rho$ affect $\left\|\mathcal{X}(\rho)\right\|$, the total content quality in FSE. Since not all values of $\rho$ induce an FSE (see Figure 2(a)), each line begins at the minimal value of $\rho$ for which an FSE exists under the corresponding parameter setting. As expected, larger values of $\rho$ encourage creators to produce more content regardless of the setting, as they get better compensation from the platform. Notably, more powerful GenAI (larger $\alpha$) reduces creator content production. Furthermore, different values of $\alpha$ align as $\rho$ approaches $1$. To see why, recall that if the platform allocates most GenAI-driven revenue to creators, creators do not suffer from the competition with GenAI.
@@
Figure 2(c) presents the platform’s optimal allocation parameter. We see that all the curves are non-increasing, implying that gradually smaller allocation parameters are enough. Focusing on the impact of high GenAI quality parameter $\alpha$ (red, green, and orange curves), we see an interesting pattern: The allocation parameter plateaus as we increase the number of creators. That is, the platform’s optimal choice tends toward higher $\rho$ regardless of ecosystem size, with larger $\alpha$ corresponding to larger optimal $\rho$. We find this surprising as Figure 2(a) indicates that more powerful GenAI permits lower values of the allocation parameter to induce an FSE. This counterintuitive finding indicates that platforms with more powerful GenAI must compensate creators more generously in terms of the allocation. For weaker GenAI, the optimal allocation parameter is similar to the minimal allocation threshold required to induce an FSE: For $\alpha=0.5$, the black curves in Figure 2(a) (minimal parameter) and Figure 2(c) (optimal parameter) are almost identical.
@@
Figures 2(d), 2(e), and 2(f) examine the platform’s revenue, creator utilities, and total quality under the platform-optimal allocation. As expected, larger creator populations increase platform revenue and total quality (increasing curves in Figure 2(d) and Figure 2(f)) while decreasing the average creator returns (decreasing curves in Figure 2(e)). Interestingly, powerful GenAI (red, green, and orange curves) yield nearly identical patterns in all graphs, suggesting that beyond a certain threshold, more powerful GenAI does not significantly impact economic outcomes. Moreover, weak GenAI (black curve) is bad for the platform compared to powerful GenAI, and good for creator utility and total content quality.
@@ 6 Discussion and Future Work
We see several promising directions for future research. First, the platform’s objective in our model is revenue maximization. However, in practical applications, the platform may have multiple objectives, such as creator retention or user welfare. Future work could explore how to balance these competing objectives. Second, while our analysis focused on proportional allocation rules, other functional forms may lead to better outcomes. For instance, allocation rules that exhibit a staircase structure, providing high revenue for certain quality levels but diminishing afterward, could potentially yield more nuanced and efficient outcomes. Third, we restricted our attention to FSEs, but other PNEs (non-FSEs) may exist. Indeed, Proposition 2 hints that an asymmetric sharing, where only some creators choose to share content, can occur in equilibrium. Understanding which allocation rules can induce such equilibria and whether they lead to improved platform revenue remains an important open question.
@@ Acknowledgements
This research was supported by the Israel Science Foundation (ISF; Grant No. 3079/24).
Appendix A Example: Prisoner’s Dilemma Dynamics
In this section, we present an example that serves a dual purpose: Illustrating our setting and notation, and revealing a Prisoner’s Dilemma-like phenomenon.
Example 7.
Consider an instance with $n=2$ creators, cost functions $c_{1}(x)=c_{2}(x)=\frac{1}{2}x^{2}$, and parameters $\mu=10$, $\alpha=\gamma=1$. Namely, the user traffic function is $T(\mathbf{x})=10\cdot\left\|\mathbf{x}\right\|$, and the GenAI quality is $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=1\cdot\mathbf{x}^{\top}\mathbf{s}$. Suppose the platform adopts the revenue allocation rule $f_{0.8}$, as defined in Equation 1 for $\rho=0.8$.
Consider the case where neither creator chooses to share data ($\mathbf{s}=(s_{1},s_{2})=(0,0)$). Since no data is available to the GenAI ($\mathbf{x}^{\top}\mathbf{s}=0$ for any $\mathbf{x}$), the GenAI’s content quality becomes $Q_{\text{AI}}(\mathbf{x},\mathbf{0})=0$. Consequently, $f_{i,\rho}(\mathbf{x},\mathbf{0})=0$ for every quality profile $\mathbf{x}$ and index $i\in[2]$; thus, each creator’s utility is the user attention directed to their content minus their production cost. In other words,
$U_{i}(\mathbf{x},\mathbf{0})=T(\mathbf{x})\frac{x_{i}}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{0})}-\frac{1}{2}x_{i}^{2}=10(x_{1}+x_{2})\frac{x_{i}}{x_{1}+x_{2}}-\frac{1}{2}x_{i}^{2}=10x_{i}-\frac{1}{2}x_{i}^{2}.$ (2)
Equation (2) suggests that utilities are decoupled: Each creator can independently maximize their utility regardless of the action of the other creator. The optimal quality strategies are attained at $x_{1}=x_{2}=10$, yielding a utility of $U_{i}((10,10),\mathbf{0})=50$ for each creator $i\in\{1,2\}$.
However, the profile $((10,10),\mathbf{0})$ is not a PNE. To see this, consider a deviation of creator 1 to full-sharing ($s_{1}^{\prime}=1$) while maintaining $x_{1}=10$. After the deviation, GenAI achieves a quality of $Q_{\text{AI}}(\mathbf{x},(1,0))=\alpha x_{1}s_{1}^{\prime}=10$. The total content quality remains $\left\|\mathbf{x}\right\|=x_{1}+x_{2}=20$, while the aggregate shared content is $\mathbf{x}^{\top}\mathbf{s}^{\prime}=x_{1}s_{1}^{\prime}+x_{2}s_{2}=10$. Furthermore, only creator 1 shares its content with GenAI; thus,
$f_{1,0.8}((10,10),(1,0))=0.8\cdot\frac{x_{1}s_{1}^{\prime}}{\mathbf{x}^{\top}\mathbf{s}^{\prime}}=0.8.$
Altogether, creator 1’s utility after the deviation is
$U_{1}^{\prime}((10,10),(1,0))$ $=T(\mathbf{x})\left(\frac{x_{1}+Q_{\text{AI}}(\mathbf{x},(1,0))\cdot f_{1,0.8}((10,10),(1,0))}{\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},(1,0))}+\right)-c_{1}(10)$
$=10\cdot 20\cdot\frac{10+10\cdot 0.8}{20+10}-\frac{1}{2}\cdot 10^{2}=70;$
thus, $U_{1}((10,10),(1,0))=70>50=U_{1}((10,10),\mathbf{0})$, implying that unilateral deviating to content sharing is strictly beneficial for creator 1, and $((10,10),\mathbf{0})$ is not a PNE.
As it turns out, the unique PNE $(\mathbf{x}^{\star},\mathbf{s}^{\star})$ under $f_{1,0.8}$ is obtained when both creators fully share their content ($\mathbf{s}^{\star}=1$) and investing symmetrically in content quality, namely $\mathbf{x}^{\star}=(9,9)$. Note that this equilibrium is an FSE and thus $\mathbf{x}^{\star}$ is an ESE. In this profile, GenAI’s quality is $Q_{\text{AI}}(\mathbf{x}^{\star},\mathbf{s}^{\star})=18$, and each creator receives a revenue allocation fraction $f_{i,0.8}=0.4$ from the AI-driven revenue. Computing utilities, we see that each creator obtains a $40.5$.
Table 1: The normal game induced by $\mathbf{x}=(10,10)$ and revenue allocation rule $f_{0.8}$.

### v2

@@ 5 Experiments
In this section, we conduct experiments to understand equilibrium outcomes and, more generally, how the allocation parameter $\rho$ impacts creator behavior. We consider cost functions of the form $c_{i}(x)=a_{i}x^{\theta}$. In the default configuration, the parameters are $n=20$, $\gamma=0.8$, $\mu=100$, $\theta=1.5$, $\alpha=0.5$, and the coefficients $\{a_{i}\}_{i=1}^{n}$ are independently drawn from the uniform distribution $Uni[1,10]$. In addition to the default configuration, we varied one parameter at a time while keeping all others fixed at their default values to assess its individual impact. Each point in Figure˜2 represents the average outcome across 150 independent game instances, with error bars indicating 95% confidence intervals obtained via bootstrapping. We discretized the possible values of $\rho$ into 100 equally-spaced points in the range $[0,1]$. For each $\rho$, we check if it induces an $\varepsilon$-FSE in $G(\rho)$ for $\varepsilon=10^{-4}$. We ran all the executions on a standard laptop, completing the entire process in a few hours. We include all implementation details in Appendix˜A.
Minimal revenue-allocation threshold
Figure˜2(a) illustrates how the minimal allocation parameter $\rho$ required for inducing an FSE varies with ecosystem parameters. All curves decrease with the number of creators, as Theorem˜5 implies: the content of each creator becomes insignificant w.r.t. all other created content, so withholding has a high alternative cost. Furthermore, more powerful GenAI (larger $\alpha$) enables lower $\rho$ choices for stable FSE, as the curves associated with a change in $\alpha$ (red, green, orange, and black curves) are ordered. The black curve $(\mu=100)$ lies above the blue curve $(\mu=50)$, suggesting that creators demand more compensation as traffic increases.
Content quality vs. allocation parameter
Figure˜2(b) shows how different values of $\rho$ affect $\left\|\mathcal{X}(\rho)\right\|$, the total content quality in FSE. Since not all values of $\rho$ induce an FSE (see Figure˜2(a)), each line begins at the minimal value of $\rho$ for which an FSE exists under the corresponding parameter setting. As expected, larger values of $\rho$ encourage creators to produce more content regardless of the setting, as they get better compensation from the platform. Notably, more powerful GenAI (larger $\alpha$) reduces creator content production. Furthermore, different values of $\alpha$ align as $\rho$ approaches $1$. To see why, recall that if the platform allocates most GenAI-driven revenue to creators, creators do not suffer from the competition with GenAI.
@@
Figure˜2(c) presents the platform’s optimal allocation parameter. We see that all the curves are non-increasing, implying that gradually smaller allocation parameters are enough. Focusing on the impact of high GenAI quality parameter $\alpha$ (red, green, and orange curves), we see an interesting pattern: the allocation parameter plateaus as we increase the number of creators. That is, the platform’s optimal choice tends toward higher $\rho$ regardless of ecosystem size, with larger $\alpha$ corresponding to larger optimal $\rho$. We find this surprising as Figure˜2(a) indicates that more powerful GenAI permits lower values of the allocation parameter to induce an FSE. This counterintuitive finding indicates that platforms with more powerful GenAI must compensate creators more generously in terms of the allocation. For weaker GenAI, the optimal allocation parameter is similar to the minimal allocation threshold required to induce an FSE: for $\alpha=0.5$, the black curves in Figure˜2(a) (minimal parameter) and Figure˜2(c) (optimal parameter) are almost identical.
@@
Figures 2(d), 2(e), and 2(f) examine the platform’s revenue, creator utilities, and total quality under the platform-optimal allocation. As expected, larger creator populations increase platform revenue and total quality (increasing curves in Figure˜2(d) and Figure˜2(f)) while decreasing the average creator returns (decreasing curves in Figure˜2(e)). Interestingly, powerful GenAI (red, green, and orange curves) yield nearly identical patterns in all graphs, suggesting that beyond a certain threshold, more powerful GenAI does not significantly impact economic outcomes. Moreover, weak GenAI (black curve) is bad for the platform compared to powerful GenAI, and good for creator utility and total content quality.
6 Extensions
In this section, we provide several natural extensions of our framework and discuss their implications on our main results.
6.1 Multiple Topics
The following extension enables creators to allocate different efforts across various content topics, while the platform maintains a single revenue-allocation rule. Consider $K$ distinct content topics. Each creator $i$ now chooses a quality vector $\mathbf{x}_{i}=(x_{i,1},\ldots,x_{i,K})\in[0,\infty)^{K}$ and a sharing vector $\mathbf{s}_{i}=(s_{i,1},\ldots,s_{i,K})\in[0,1]^{K}$, where $x_{i,k}$ and $s_{i,k}$ represents their quality and sharing level in topic $k$, respectively. The cost function, which now gets $\mathbf{x}_{i}$ as input, is convex but need not be separable across topics, i.e., $c_{i}(\mathbf{x}_{i})=\sum_{k}c_{i,k}(x_{i,k})$, for most results to hold.
Each topic $k$ generates independent user traffic $T_{k}$, and the GenAI system’s quality is $Q_{\text{AI},k}$; both depend only on the restriction of $\mathbf{x}$ and $\mathbf{s}$ to topic $k$ and have the same functional forms as in the single-topic case. The platform commits to a single allocation parameter $\rho\in[0,1]$ across all topics. Under proportional allocation, creator $i$’s revenue from topic $k$ is given by
$R_{i,k}(\mathbf{x},\mathbf{s};f_{\rho})=\frac{T_{k}(\mathbf{x})x_{i,k}}{\sum_{j=1}^{n}x_{j,k}+Q_{\text{AI},k}}+\frac{T_{k}(\mathbf{x})Q_{\text{AI},k}}{\sum_{j=1}^{n}x_{j,k}+Q_{\text{AI},k}}\cdot\frac{\rho x_{i,k}s_{i,k}}{\sum_{j=1}^{n}x_{j,k}s_{j,k}},$
and its overall utility is $U_{i}(\mathbf{x},\mathbf{s};f_{\rho})=\sum_{k=1}^{K}R_{i,k}(\mathbf{x},\mathbf{s};f_{\rho})-c_{i}(\mathbf{x}_{i})$.
Results
Lemma˜2 extends to each topic independently: for fixed quality choices, the optimal sharing decision $s_{i,k}$ for topic $k$ follows the same threshold structure in LABEL:{eq:sharing-threshold-form}, with $\tau_{i,k}=\frac{x_{i,k}}{x_{i,k}+\sum_{j\neq i}x_{j,k}+\alpha_{k}\sum_{j\neq i}x_{j,k}s_{j,k}}$. Lemma˜3 holds as well—since the only interplay between different topics is through the cost function, the same analysis still applies. Theorem˜5 requires some modifications, but the general idea remains the same: a large enough $\rho$ admits FSE. The main technical difference is establishing per-topic rational quality bounds $\mathbf{x}_{i,k}^{\max}$ when the cost functions couple topics together. The approach of Algorithm˜1 generalizes to this setting too, but testing the condition in Line 7 becomes computationally intensive, as we need to check each of the $2^{K}-1$ possible sharing deviations. This can be avoided when the cost functions are separable across topics.
6.2 Prior GenAI Data
Next, we allow the GenAI system to possess a fixed body of prior data, independent of the creators’ sharing contributions. Formally, we introduce a non-strategic creator indexed by 0 with fixed quality $x_{0}>0$ and full sharing level $s_{0}=1$. The quality of the GenAI system then becomes $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=\alpha\left(x_{0}+\sum_{j=1}^{n}x_{j}s_{j}\right)$, so that even if no strategic creator shares content, the GenAI still attains positive quality $\alpha x_{0}$. This modification also affects user traffic. Essentially, we assume that there is another source of $x_{0}$ body of content in the platform that is always available to the GenAI system.
Results
From creator $i$’s perspective, the new creator is just another content contributor with fixed quality and full sharing; thus, the threshold from Lemma˜2 updates accordingly to
$\tau_{i}=\frac{x_{i}}{x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}+(1+\alpha)x_{0}}.$
Notably, prior data $x_{0}>0$ strictly decreases $\tau_{i}$, making full sharing more attractive for creators. The proof of Lemma˜3 requires a minor modification, as the dummy creator’s strategy is fixed. Nonetheless, the resulting subgame among the creators maintains the DSC property. All other results extend directly to this setting.
6.3 Alternative Optimization Objectives
While our analysis in Section˜4 focused on maximizing platform revenue (recall (PO)), there are other desirable objectives to consider. For instance, a platform may wish to balance its retained value with total content production and creator welfare. Our framework naturally extends to such cases, where the target function is Lipschitz continuous in the quality profile $\mathbf{x}$ and the allocation parameter $\rho$. The platform’s problem becomes
$\sup_{\rho\in[0,1]}\ f(\mathcal{X}(\rho),\rho)\quad\text{subject to }(\mathcal{X}(\rho),\mathbf{1})\text{ is an FSE}.$
Natural choices for $f$ include total content production $f(\mathbf{x},\rho)=\|\mathbf{x}\|_{1}$, creator welfare $f(\mathbf{x},\rho)=\sum_{i=1}^{n}U_{i}(\mathbf{x},\mathbf{1};f_{\rho})$, or regularized combinations such as
$f(\mathbf{x},\rho)=(1-\lambda)U_{P}(\mathbf{x},\mathbf{1};f_{\rho})+\lambda\|\mathbf{x}\|_{1}.$
Results
The same algorithmic approach from Algorithm˜1 applies directly to these generalized objectives, with the approximation guarantees of Theorem˜8 carrying over after adjusting the constants $A$ and $B$ to account for the Lipschitz properties of $f$.
6.4 GenAI-Dependent Traffic
In our main model, user traffic $T(\mathbf{x})$ depends solely on the quality of human-generated content. However, high-quality AI-generated content may also attract users to the platform. To capture this effect, we extend our traffic model to incorporate both human and AI contributions:
$T(\mathbf{x},\mathbf{s})=\mu\left(\left\|\mathbf{x}\right\|+Q_{\text{AI}}(\mathbf{x},\mathbf{s})\right)^{\gamma}.$ (3)
All other components of the model remain unchanged.
Results
Under this extended traffic model, the creator’s optimal sharing decision retains its threshold structure, with an adjusted threshold given by $\tau_{i}=\frac{1}{\alpha}\left[\left(\frac{\sum_{j}x_{j}+\alpha\sum_{j}x_{j}s_{j}}{\sum_{j}x_{j}+\alpha\sum_{j\neq i}x_{j}s_{j}}\right)^{1-\gamma}-1\right]$. Notably, larger values of $\gamma$ decrease $\tau_{i}$, making sharing more attractive. This occurs because higher traffic elasticity amplifies the positive externality that sharing exerts on the platform’s traffic.
The analysis of ESE remains tractable under this extension. In the enforced sharing game where $\mathbf{s}=\mathbf{1}$, the traffic function simplifies to $T(\mathbf{x},\mathbf{1})=\mu(1+\alpha)^{\gamma}\left\|\mathbf{x}\right\|^{\gamma}$. This expression is structurally identical to our original traffic model, differing only by a multiplicative scaling of the baseline parameter $\mu$ by $(1+\alpha)^{\gamma}$. Consequently, Lemma˜3 applies directly: the subgame $G(\rho)$ under enforced sharing admits a unique ESE. Since the utility function at $\mathbf{s}=\mathbf{1}$ retains its structural properties, our main results carry over to this setting with adjusted constants. Specifically, the FSE existence condition (Theorem˜5) holds by redefining $x_{i}^{\max}$ as the unique positive solution to $c_{i}(x)=\mu(1+\alpha)^{\gamma}x^{\gamma}$. The approximation guarantees of Theorem˜8 follow similarly.
6.5 GenAI Quality Beyond Linearity
In our baseline model, the quality of the GenAI system is linear in the aggregate amount of shared content, given by $Q_{\text{AI}}(\mathbf{x},\mathbf{s})=\alpha\sum_{i=1}^{n}x_{i}s_{i}$. This specification is both analytically convenient and captures the essence of the strategic dynamics in the system. However, in practice, GenAI systems often exhibit diminishing marginal returns as the amount of training data grows [31]. To capture this effect, we consider the following generalization of $Q_{\text{AI}}$:
$Q_{\text{AI}}(\mathbf{x},\mathbf{s})=\alpha\left(\sum_{i=1}^{n}x_{i}s_{i}\right)^{\beta},\qquad\beta\in[0,1].$
This form, introduced by Yao et al. [50], captures diminishing returns in GenAI quality. Notice that when $\beta=1$, this formulation reduces to our original model.
Theoretical results.
When $\beta<1$, the enforced sharing game is no longer concave and therefore does not satisfy Rosen’s DSC condition. As a consequence, we lose general guarantees on the existence and uniqueness of an ESE. Indeed, in numerical experiments, we observe instances that admit multiple approximate ESEs, which can differ in the induced quality profiles. From a more positive perspective, several qualitative insights from our baseline model remain valid. First, Proposition˜4 transfers directly: when $\rho=1$, full sharing is a dominant strategy for every creator, and the subgame admits a unique PNE that is an FSE. Second, a weaker analogue of Lemma˜2 continues to hold. In particular, when no creator shares content, creator $i$ has an incentive to deviate and share whenever $\rho>\frac{x_{i}}{\sum_{j=1}^{n}x_{j}}$. This mirrors the baseline threshold and gives rise to a prisoner’s dilemma-like structure in large systems.
Simulation methodology.
To study equilibria under the generalized GenAI quality function with $\beta<1$, we rely on numerical methods. We focus on the enforced sharing game, where all creators are required to share ($\mathbf{s}=\mathbf{1}$), and compute an approximate ESE via best-response dynamics. Starting from an arbitrary initial quality profile, creators update their quality choices sequentially. Each best response in $x_{i}$ is computed via exact one-dimensional optimization over the interval $0,x_{i}^{\max}]$, where $x_{i}^{\max}$ is the analytical upper bound derived in the paper. To ensure robust convergence, we employ a damping factor that interpolates between the current quality and the computed best response. We declare convergence once the maximum change in any creator’s quality falls below a tolerance $\varepsilon$. Across all instances we tested, these dynamics converged quickly to a stable quality profile, which we treat as an $\varepsilon$-ESE.
To test whether a converged profile constitutes an $\varepsilon$-FSE, we explicitly check unilateral deviations in the sharing decision. For each creator, we consider both $s_{i}=0$ and $s_{i}=1$, and for each case, compute the optimal quality choice via one-dimensional optimization. If no such deviation yields a utility gain exceeding $\varepsilon$, we classify the profile as an $\varepsilon$-FSE. Notably, although multiple approximate equilibria may exist in this setting, our simulations reveal that they exhibit similar qualitative properties across different initializations.
$\alpha=0.5$
(a) Min. $\rho$ inducing $\varepsilon$-FSE
(b) $\rho$ vs. $\left\|\mathcal{X}(\rho)\right\|$ in FSE
(c) Optimal allocation parameter $\rho$
$\alpha=5$
(d) Min. $\rho$ inducing $\varepsilon$-FSE
(e) $\rho$ vs. $\left\|\mathcal{X}(\rho)\right\|$ in FSE
(f) Optimal allocation parameter $\rho$
Figure 3: Effect of the diminishing returns parameter $\beta$ on equilibrium outcomes for $\alpha=0.5$ (top row) and $\alpha=5$ (bottom row).
Simulation results.
We examine the effect of $\beta\in\{0.2,0.4,0.6,0.8,1.0\}$ on equilibrium outcomes, using two values of the GenAI efficiency parameter: $\alpha=0.5$ (moderate GenAI) and $\alpha=5$ (powerful GenAI). The results are presented in Figure˜3.
For moderate GenAI ($\alpha=0.5$), several baseline trends persist. The minimal allocation parameter required to induce an FSE decreases with the number of creators, consistent with the intuition from Theorem˜5. Content quality continues to increase with $\rho$. However, we observe that the minimal $\rho$ increases with $\beta$, which is somewhat counterintuitive: in our baseline analysis, stronger GenAI (larger $\alpha$) decreased the minimal $\rho$, yet here a different notion of “stronger” GenAI (larger $\beta$, corresponding to better data utilization) has the opposite effect. The optimal allocation parameter in this regime closely tracks the minimal threshold, suggesting that the platform retains as much revenue as possible while maintaining an FSE.
For powerful GenAI ($\alpha=5$), the relationship between $\beta$ and the minimal $\rho$ becomes non-monotonic, though the decrease with the number of creators persists. A clearer pattern emerges in the optimal allocation: larger $\beta$ leads to higher optimal $\rho$, and the gap between optimal and minimal allocation parameters widens substantially for larger $\beta$. This indicates that with powerful GenAI exhibiting near-linear returns, the platform benefits from more generous revenue sharing to incentivize content production. The content-quality curves confirm that larger $\beta$ reduces total content at any given $\rho$, though quality remains increasing in the allocation parameter.
Overall, the core tradeoffs identified in our baseline model–balancing revenue retention against content production incentives–persist across the range of $\beta$ values. However, the interaction between $\alpha$ and $\beta$ introduces additional complexity, and some of the clean monotonicity properties from the linear case are lost when both parameters vary.
@@ 7 Discussion and Future Work
We see several promising directions for future research. First, while our analysis focused on proportional allocation rules, other functional forms may lead to better outcomes. For instance, allocation rules that exhibit a staircase structure, providing high revenue for certain quality levels but diminishing afterward, could potentially yield more nuanced and efficient outcomes. Second, we restricted our attention to FSEs, but other PNEs (non-FSEs) may exist. Indeed, Lemma˜2 hints that asymmetric sharing, where only some creators choose to share content, can occur in equilibrium. Understanding which allocation rules can induce such equilibria and whether they lead to improved platform revenue remains an important open question. Third, extending our theoretical guarantees to more general GenAI quality functions with diminishing returns ($\beta<1$ in Section˜6.5) remains open. The loss of concavity invalidates our uniqueness results, and while simulations suggest equilibria persist with similar qualitative behavior, formal existence and characterization guarantees are lacking. Finally, our model considers a static interaction; incorporating dynamic elements such as creators entering or exiting the platform, evolving GenAI capabilities, or repeated interactions with learning could yield richer insights into long-term platform-creator relationships.
@@ Acknowledgements
This research was supported by the Israel Science Foundation (ISF; Grant No. 3079/24). We thank the anonymous reviewers for their valuable feedback.
Appendix A Omitted Simulation Details
This section provides technical details about the algorithms and computational resources used in our experimental simulations. [^4] Our code is available at https://github.com/GurKeinan/Code_for_Strategic_Content_Creation_with_GenAI_To_Share_or_Not_to_Share_Paper.
A.1 ESE Computation
Theoretical algorithm.
Algorithm 2 provides a theoretically justified method for approximating the unique ESE for a given revenue-allocation parameter $\rho$. It employs multi-agent mirror descent with gradient-based updates to iteratively refine each creator’s quality choice $x_{i}$ while maintaining non-negativity constraints. The strong monotonicity of the enforced sharing game (established in Lemma˜3) guarantees convergence at a rate of $O(1/T)$ [12], where $T$ is the number of iterations.
Algorithm 2 Multi-Agent Mirror Descent for ESE Quality Profile (adapted from Bravo et al. [12])
2:Approximate ESE quality profile $x^{(T)}$.
3:Initialize quality profile $x\leftarrow x^{(0)}$
4:for $t\leftarrow 0$ to T1T-1 do
5:  Calculate gradient vector $v^{(t)}\in\mathbb{R}^{n}$. For each creator $i$:
6:  for $i\leftarrow 1$ to $n$ do
7:   Let $X^{(t)}=\sum_{j=1}^{n}x_{j}^{(t)}$ and $X_{-i}^{(t)}=X^{(t)}-x_{i}^{(t)}$
8:   $v_{i}^{(t)}\leftarrow\frac{\mu(1+\alpha\rho)}{1+\alpha}(X^{(t)})^{\gamma-2}\,[\gamma x_{i}^{(t)}+X_{-i}^{(t)}]-c_{i}^{\prime}(x_{i}^{(t)})$
9:  Update candidate profile: $x_{cand}\leftarrow x^{(t)}+\eta_{t}v^{(t)}$
10:  Project onto feasible set: $x^{(t+1)}\leftarrow\max(0,x_{cand})$
11:return $x^{(T)}$
Practical implementation.
While Algorithm 2 provides provable convergence guarantees, we adopted a more practical approach in our simulations. Since each creator’s utility is strictly concave in their own quality $x_{i}$ (as established in Lemma˜3), the ESE is uniquely characterized by the first-order conditions: for each creator $i\in[n]$, the equilibrium quality satisfies